Lane Industries is considering three independent projects, each of which requires a $3.4 million investment. The estimated internal rate of return (IRR) and cost of capital for these projects are presented here:

Project H (high risk): Cost of capital = 13% IRR = 15%
Project M (medium risk): Cost of capital = 11% IRR = 9%
Project L (low risk): Cost of capital = 10% IRR = 11%

Note that the projects' costs of capital vary because the projects have different levels of risk. The company's optimal capital structure calls for 40% debt and 60% common equity, and it expects to have net income of $4,200,000. If Lane establishes its dividends from the residual dividend model, what will be its payout ratio? Round your answer to 2 decimal places.
%
Answer #1

Acceptable projects are those whose IRR is greater than the cost of capital

Hence, project H and L are acceptable

Required investment = 3,400,000*2 =$6,800,000

To be funded from equity = 60%*6,800,000 =$4,080,000

Under residual dividend model, the earnings remaining after meeting all profitable projects is distributed as dividend

Payout ratio = dividend/net income

= (4,200,000-4,080,000)/4,200,000

=2.8571%

I.e. 2.86%